# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

### v1

@@ 3.1. Case A ($a=0$)
    Therefore, using Maple, we get
    $l=\frac{c}{32(c^{2}-1)}<0$
    and we may conclude that the limit cycle is stable for $10<c<$ 1 and the bifurcation is subcritical – [27, Remarks 1 and 2, pp. 383]. In this case $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable for $\mu>0$ with an asymptotically stable periodic solution for $\mu<0$. The same conclusion holds for $b$.
@@
    Using linear analysis, one knows that Hopf bifurcation exists only if $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0$ and $\mbox{Det}(E_{2}),\mbox{Det}(E_{3})>0.$ Indeed,
    $1\displaystyle 3{c}^{2}-{b}^{2}-2{c}^{2}b=0\Leftrightarrow b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>$ 1 .
    Figure 3. The bifurcation curves associated to (3.1) for Case A. Within each depicted region, phase portraits are $C^{1}$–equivalent. (a) The global phase portraits on the Poincaré disc corresponding to Regions a–p have been sketched in Figures 8. (b) The local portraits corresponding to Regions 1–28 have been plotted in Table 1 of Appendix.
@@
    Let $b=c\left(-c\pm\sqrt{c^{2}+3}\right).$ The Jacobian matrix at $E_{2,3}$ has a pair of purely imaginary eigenvalues ±$\pm i\sqrt{2c\sqrt{c^{2}+3}-2-2c^{2}}$ with $c>1.$ Let $b=-c^{2}\pm c\sqrt{c^{2}+3}+\mu$ where $\mu\in\mathbb{R}$ denotes the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ associated with the Jacobian matrix at $E_{2,3}$
@@
    There is a pair of complex conjugate eigenvalues when $\sigma<0$ and
@@
    Since the first Lyapunov coefficient is
    $l={\frac{c\left(\left(4c^{2}+3\right)\left(\sqrt{c^{2}+3}-c\right)-3c\right)\left(\sqrt{c^{2}+3}c\left(21+44c^{2}+16c^{4}\right)-c^{2}\left(69+68c^{2}+16c^{4}\right)-9\right)}{32\left(\sqrt{c^{2}+3}c-c^{2}-1\right)^{2}\left(\sqrt{c^{2}+3}-c\right)^{7}}},$
    we may deduce, using Maple, that these limit cycles are unstable for $c>0$.
@@
Figure 4. The classification of the equilibrium point $E_{1}$ and the corresponding bifurcation curves for (3.1) in Case B. In this figure, un, sn, uf and sf refer to unstable (stable) node and unstable (stable) focus, respectively. (a) The local phase portraits associated with Regions 1–8 have been plotted in Figure 2. (b) The global phase portraits on the Poincaré disc associated with Regions I–IV have been plotted in Figure 9.
@@
Numerically, we may observe that the non-degenerate Hopf bifurcation of items (3) and (5) of Theorem 3.4 are, respectively, subcritical and supercritical. The bifurcation $T_{H}^{1}$ yields a stable periodic solution, say $C^{s}$, and $T_{H}^{2,3}$ generates two unstable and $\kappa$–symmetric periodic solutions – confirm on Table 1 of Appendix.
@@ 3.2. Case B ($b=0$)
With respect to the vector field $f_{(a,0,c)}$ of (3.1), the global representations of the transition/bifurcation curves in the space of parameters $\0(a,c)\in\mathbb{R}\times\mathbb{R}\backslash\{0\}$ are as follows:
@@
    The Hopf bifurcation is supercritical at $a=-1$ and subcritical at $a=1$.
Schematic bifurcation lines for $b=0$ have been plotted in Figure 4. Illustrative pictures in the phase space have been depicted in Figure 9 in the case where the eigenvalues of (3.5) at $E_{1}$ are complex non-real.
@@
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues ±$\pm i.$ Let $a=\pm 1+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix (3.5) at $E_{1}$ are (computations performed with Maple):
@@
    For $\mu$ close to zero, we have $04{c}^{2}{\mu}^{2}+4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4<$ 0, which means that $\lambda(\mu)$ is complex non-real. Furthermore $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}\big{|}_{\mu=0}=\mp c\neq 0.$ Following [26, Equations (2) and (3)], the first Lyapunov coeficient is $l=-\frac{1}{32}c$ which is negative for $c>0.$ Hence the bifurcation is subcritical at $a=1$ and supercritical for $a=-1$.
    For the case $a=-1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable focus for $\mu>0$ with an asymptotically stable periodic solution emerging for $\mu>0$. For the case $a=1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ and unstable focus for $\mu<0$ with an asymptotically stable periodic solution emerging $\mu>0$. This finishes the proof.
Remark 3.9.
We may observe in Figure 9 that, for $c>0$ (resp. $c<0$), the Hopf bifurcation (with respect to $a$) is supercritical for $a=-1$ giving rise to a stable (resp. unstable) periodic solution which disappears at a subcritical Hopf bifurcation at $a=1,$ see regions 5-8 of Table 2.
3.3. Case C ($a>0,$ $10<b<$ 1)
@@
For $10<b<$ 1, the discriminant satisfies $\Delta<0$. According to Cardano’s formula, this implies that the polynomial $H(x)$ has exactly one real root, denoted by $x^{\star}.$ Then the equilibrium of (3.1) is explicitly given by
@@
The following assertions hold for the vector field $f_{(a,b,c)}$ of (3.1), under the conditions $a>0,$ and $10<b<$ 1:
@@
    A non-degenerate Hopf bifurcation occurs at $c^{2}=\frac{b}{1-{x^{*}}^{2}}.$ The resulting limit cycle is attracting when $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and repelling for $c<\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
Note that $x^{\star}\neq\pm 1$ (if $x^{\star}=\pm 1$, then $b=0$, which is a contradiction). Illustrative pictures in the phase space have been depicted in Figure 10. As before, we use the notation $\mbox{Tr}(E_{1})$ and $\mbox{Det}(E_{1})$ the trace and determinant operators of $Df_{(a,b,c)},$ evaluated at $E_{1}$.
@@
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues ±$\pm i\sqrt{b{x^{*}}^{2}-b+1}.$ Let $b=c^{2}\left(1-{x^{*}}^{2}\right)+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix at $E_{1},$
@@
The statement of Proposition 3.10 says nothing about the criticality of the Hopf Bifurcation. Analytically, it is very difficult to compute the first Lyapunov exponent of (3.1) at an equilibrium whose explicit expression we do not know. However, numerics of Figure 10 suggest that the Hopf bifurcation is supercritical for for $c=\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and subcritical for $c=-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
4. Canards for the singular case $c\to+\infty$ in Cases A and B
The goal of this section is the study of the dynamics of (3.1) when $|c|\to+\infty$. Many results concerning the asymptotic behaviour of the solution of the FN model as $|c|\to+\infty$ are available in the literature – see [21, 22, 23]. Our concern is to match these results with the local bifurcation study performed in Section 3. We are particularly interested in Cases A and B whose results have been described in Theorems 3.4 and 3.8. By changing the time scale of (3.1), considering $\tau\to t/c$ and $\varepsilon\to 1/c^{2},$ we obtain an equivalent version of (3.1) as follows (written with respect to the slow time $\tau$):
$\left\{\begin{array}[]{l}\varepsilon\dot{x}=\displaystyle\left[y-\left(\frac{x^{3}}{3}-x\right)\right]=:f(x,y,\varepsilon)\\ \\ \dot{y}=\displaystyle-(x-a+by)=:g(x,y,\varepsilon)\end{array}\right.$ (4.1)
where $a,b\in\mathbb{R}$ and $\0c\in\mathbb{R}\backslash\{0\}.$ For $0<\varepsilon\ll 1,$ or equivalently $|c|\gg 1,$ system (4.1) is what we call a fast-slow system (see [10, 14] and references therein).
A fast-slow system is a system of differential equations in which some variables have their derivatives with larger magnitude than others. This leads to a two time-scale system. The general approach to this type of system starts by grouping the variables in two disjoint sets: fast variable $x$ and slow variable $y$. By allowing the dynamics to be separated into two distinct phases, slow and fast, we can study each one independently and understand its underlying dynamics. This separation is introduced in system (4.1) via the parameter $\varepsilon.$
@@ 4.1. The critical manifold and bifurcations
$\mathcal{C}_{0}=\{(x,y)\in\mathbb{R}^{2}:f(x,y,0)=0\}=\{(x,y)\in\mathbb{R}^{2}:y=x^{3}/3-x\}$ (4.2)
@@
    The subset $S\subset\mathcal{C}_{0}$ is said to be normally hyperbolic if for all $p=(x,y)\in{S}$, we have:
[…]
@@
    A normally hyperbolic set $S$ is said to be attracting (resp. repelling) if, for all $p\in{S}$, all eigenvalues of $\frac{\partial f}{\partial x}\left(p,0\right)$ have negative (resp. positive) real part, respectively.
@@
With respect to (4.1) with critical manifold $\mathcal{C}_{0}$, the point $p\in\mathcal{C}_{0}$ is said to be a fold point if:
$\frac{\partial f}{\partial x}\left(p,0\right)=0\,,\quad\frac{\partial^{2}f}{\partial x^{2}}\left(p,0\right)\neq 0\quad\text{ and }\quad\frac{\partial f}{\partial y}\left(p,0\right)\neq 0.$
If $g\left(p,0\right)\neq 0$, the fold point is called regular.
Coming back to system (4.1), we have $f\left(x,y,0\right)=y-x^{3}/3+x$ and, for $p=(x,y)\in\mathcal{C}_{0}$, one has:
$\frac{\partial f}{\partial x}\left(p,0\right)=x^{2}-1\neq 0\qquad\text{ if }\,\,x\neq\pm 1.$
Therefore, if $b=3/2$, the equilibria $E_{2}=(-1,2/3)$ and $E_{3}=(1,-2/3)$ are fold points (according to the Definition 4.2) since:
@@
and $S_{0}=\mathcal{C}_{0}\setminus\left\{E_{2},E_{3}\right\}$ is normally hyperbolic. We may split $S_{0}$ into three disjoint open subsets:
@@
where $\mathcal{C}_{0\text{L}}$ and $\mathcal{C}_{0\text{R}}$ are repelling subsets and $\mathcal{C}_{0\text{M}}$ is attracting. This means that in the fast flow, trajectories of (4.1) move backwards either to $\mathcal{C}_{0\text{L}}$ or $\mathcal{C}_{0\text{R}}$ and towards to $\mathcal{C}_{0\text{M}}$ – see Figure 5.
@@
Yet in the singular case ($\varepsilon=0$), depending on the values of $a,b\in\mathbb{R}$, there exists at least one and at most three equilibria, as result of the intersection of the cubic curve, $f(x,y,0)=0,$ with the line defined by $x-a+by=0$ ($\Leftrightarrow\dot{y}=0).$ Moreover, when the system has three equilibria only one of the following scenarios may occur (cf. [10]):
@@
either all three equilibria belong to $\mathcal{C}_{0\text{M}}$ or else
@@
they are each one in a distinct region, $\mathcal{C}_{0\text{L}},$ $\mathcal{C}_{0\text{M}}$ and $\mathcal{C}_{0\text{R}}$.
The independence of the time scales when $\varepsilon=0$ makes the system simpler to understand. Fenichel’s Theorem [6] says that, near the normally hyperbolic part of $\mathcal{C}_{0}$, when $0<\varepsilon\ll 1,$ the dynamics of the perturbed system is similar to that of the singular system, with a deviation of $\mathcal{O}\left(\varepsilon\right).$ The smaller the $\varepsilon,$ the more similar the system trajectories are to those described in the singular system ($\varepsilon=0$). Fenichel’s Theorem guarantees that for $0<\varepsilon\ll 1,$ there exists a set $S_{\varepsilon}$, $C^{1}$–close to $S_{0}$ (in the Haussdorf topology), that exhibits the same behaviour as $S_{0}$.
@@
The set $S_{\varepsilon}$ is referred to as the slow manifold of system (4.1).
When the system’s equilibrium is also a fold point, a supercritical (with respect to $b$) Hopf bifurcation occurs, giving rise to a periodic solution. In this section, the canard phenomenon described in [12, 13] bridges the local periodic solution from the Hopf bifurcation to the global limit cycle, as we proceed to explain.
@@
A trajectory segment of a fast-slow system (4.1) is a canard if it stays within $\mathcal{O}\left(\varepsilon\right)$ distance to a repelling branch of a slow manifold for a time that is $\mathcal{O}\left(1\right)$ on the slow time scale $\tau=t\varepsilon$.
Canards are related to equilibria of (4.1) that occur at fold points of the slow manifold. The next definitions deal with the general equation parametrised by $\lambda\in\mathbb{R}$:
@@
We now introduce the definition of singular fold point, which differs from that of a fold point – it is required to coincide with an equilibrium point of (4.1):
@@
With respect to (4.4), let $p=\left(x,y\right)$ be a fold point. The point $p$ is called a singular fold if:
$f\left(p,\lambda,0\right)=0,$ $\frac{\partial f}{\partial x}\left(p,\lambda,0\right)=0,$ (4.5)
$\frac{\partial^{2}f}{\partial x^{2}}\left(p,\lambda,0\right)\neq 0,$ $\frac{\partial f}{\partial y}\left(p,\lambda,0\right)\neq 0\quad\text{and}\quad g\left(p,\lambda,0\right)=0.$
@@
A singular fold point $p$ is said to be regular if:
$\frac{\partial g}{\partial x}\left(p,\lambda,0\right)\neq 0\quad\text{ and }\quad\frac{\partial g}{\partial\lambda}\left(p,\lambda,0\right)\neq 0.$ (4.6)
The behaviour around a regular singular fold point is described in the next theorem. The existence of canards is coupled with the existence of a Hopf bifurcation, which in such systems is referred to as a singular Hopf bifurcation.
@@
Consider a fast-slow system of the form (4.4) where $\left(x,y\right)=\left(0,0\right)$ is a regular singular fold point for $\lambda=0$, and that may be written in the following form:
@@
Assume that, for $\varepsilon=0$, there is a slow trajectory connecting the repelling and attracting regions of the critical manifold $\mathcal{C}_{0}$. Then, there exist $\varepsilon_{0}>0$ and $\lambda_{0}>0$ such that for $0<\varepsilon<\varepsilon_{0}$ and $|\lambda|<\lambda_{0}$, the system has an equilibrium point $p\in\mathbb{R}^{2}$ near the origin where $p\to\left(0,0\right)$ as $\left(\lambda,\varepsilon\right)\to\left(0,0\right)$. Then, there exists a smooth function $\lambda_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a value $\lambda$ that gives rise to a family of canards, asymptotically defined by:
$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\left(B+A\right)\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$
and there exists a continuous function $\lambda_{H}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each value of $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $\lambda$ of Hopf bifurcations in the system, asymptotically defined by:
$\lambda_{H}\left(\sqrt{\varepsilon}\right)=-B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$

### v3

@@ 3.1. Case A ($a=0$)
    Using Maple, we get
    $\ell(c)=\frac{c}{32(c^{2}-1)}<0$
    and we may conclude that the limit cycle is stable for $0<c<1$ and the bifurcation is supercritical – [27, Remark 1, pp. 383]. In this case, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ ($\Leftrightarrow b>c^{2}$) and unstable for $\mu<0$ ($\Leftrightarrow b<c^{2}$), with an asymptotically stable periodic solution for $\mu<0$ ($\Leftrightarrow b<c^{2}$).
@@
    Using linear analysis, one knows that Hopf bifurcation exists only if $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0$ and $\mbox{Det}(E_{2}),\mbox{Det}(E_{3})>0.$ Indeed, we have:
    $\mbox{Tr}(E_{2})=\mbox{Tr}(E_{3})=0\mbox{ and }b>1$
    $\Leftrightarrow$ $3{c}^{2}-{b}^{2}-2{c}^{2}b=0\mbox{ and }b>1$
    $\Leftrightarrow$ $b=c\left(-c\pm\sqrt{c^{2}+3}\right)\mbox{ and }b>1.$
    Figure 3. The bifurcation curves associated to (3.1) for Case A. The local portraits corresponding to Regions 1–28 have been plotted in Table 1 of Appendix.
@@
    Let $b=c\left(-c\pm\sqrt{c^{2}+3}\right).$ The Jacobian matrix at $E_{2,3}$ has a pair of purely imaginary eigenvalues $\pm i\sqrt{2c\sqrt{c^{2}+3}-2-2c^{2}}$ with $c>1.$ Let $b=-c^{2}\pm c\sqrt{c^{2}+3}+\mu$ where $\mu\in\mathbb{R}$ denotes the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ associated with the Jacobian matrix at $E_{2,3}$
@@
    There is a pair of complex conjugate eigenvalues when $\sigma<0$. In addition, we have:
@@
    The first Lyapunov coefficient $\ell$ associated to $f_{(0,b,c)}$ at $E_{2,3}$ depends on $c$ and is given explicitly by (computations in Maple):
    $\ell(c)={\frac{c\left(\left(4c^{2}+3\right)\left(\sqrt{c^{2}+3}-c\right)-3c\right)\left(\sqrt{c^{2}+3}c\left(21+44c^{2}+16c^{4}\right)-c^{2}\left(69+68c^{2}+16c^{4}\right)-9\right)}{32\left(\sqrt{c^{2}+3}c-c^{2}-1\right)^{2}\left(\sqrt{c^{2}+3}-c\right)^{7}}},$
    Since $\lim_{c\to 1^{+}}\ell(c)=\lim_{c\to+\infty}\ell(c)=+\infty$ and $\frac{\partial\ell(c)}{\partial c}=0\quad\text{at}\quad c^{\star}\approx 1.315$, then $\ell$ remains positive in the neighbourhood of $c^{\star}$. Therefore, the corresponding limit cycles for $c>1$ are unstable and the bifurcation is subcritical.
@@
Numerically, we may observe that the non-degenerate Hopf bifurcation of items (3) and (5) of Theorem 3.4 are, respectively, supercritical and subcritical. The bifurcation $T_{H}^{1}$ yields a stable periodic solution, say $C^{s}$, and $T_{H}^{2,3}$ generates two unstable and $\kappa$–symmetric periodic solutions – confirm on Table 1 of Appendix.
@@ 3.2. Case B ($b=0$)
[…]
@@
    Figure 4. The classification of the equilibrium point $E_{1}$ and the corresponding bifurcation curves for (3.1) in Case B. In this figure, un, sn, uf and sf refer to unstable (stable) node and unstable (stable) focus, respectively. The local phase portraits associated with Regions 1–8 have been plotted in Table 2.
@@
    The Hopf bifurcation is supercritical(resp. subcritical) if $c>0$ (resp. $c<0$).
Schematic bifurcation lines for $b=0$ have been plotted in Figure 4. The local phase portraits associated with Regions 1–8 have been plotted in Table 2
@@
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues $\pm i.$ Let $a=\pm 1+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix (3.5) at $E_{1}$ are (computations performed with Maple):
@@
    For $\mu$ close to zero, we have $4{c}^{2}{\mu}^{2}+4{c}^{2}{\mu}^{3}+{c}^{2}{\mu}^{4}-4<0$, which means that $\lambda(\mu)$ is complex non-real. Furthermore
    $\frac{d\,{\rm Re}{\lambda(\mu)}}{d\mu}\big|_{\mu=0}=\mp c\neq 0.$
    The first Lyapunov coefficient depends on $c$ and is given by
    $\ell(c)=-\frac{1}{32}c$
    which has opposite sign to that of $c\neq 0$. Hence the bifurcation is subcritical for $c<0$ and supercritical for $c>0$.
    For the case a=1a=-1, $E_{1}$ is an asymptotically stable equilibrium for $\mu<0$ and unstable focus for $\mu>0$ with an asymptotically stable periodic solution emerging for $\mu>0$. For the case $a=1$, $E_{1}$ is an asymptotically stable equilibrium for $\mu>0$ and unstable focus for $\mu<0$ with an asymptotically stable periodic solution emerging $\mu>0$. This finishes the proof.
3.3. Case C ($a>0,$ $0<b<1$)
@@
For $0<b<1$, the discriminant satisfies $\Delta<0$. According to Cardano’s formula, this implies that the polynomial $H(x)$ has exactly one real root, denoted by $x^{\star}.$ Then the equilibrium for system (3.1) is explicitly given by
@@
The following assertions hold for the vector field $f_{(a,b,c)}$ of (3.1), under the conditions $a>0,$ and $0<b<1$:
@@
    A non-degenerate Hopf bifurcation occurs at $c^{2}=\frac{b}{1-{x^{*}}^{2}}.$ The numerical simulations indicate the resulting limit cycle is attracting when $c>\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and repelling for $c<-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
Note that $x^{\star}\neq\pm 1$ (if $x^{\star}=\pm 1$ then $b=0$, which is a contradiction).
As before, we use the notation $\mbox{Det}(E_{1})$ and $\mbox{Tr}(E_{1})$ for the determinant and trace operators of $Df_{(a,b,c)},$ evaluated at $E_{1}$.
@@
    The Jacobian matrix at $E_{1}$ has a pair of purely imaginary eigenvalues $\pm i\sqrt{b{x^{*}}^{2}-b+1}.$ Let $b=c^{2}\left(1-{x^{*}}^{2}\right)+\mu$ where $\mu\in\mathbb{R}$ refers to the bifurcation parameter. The eigenvalues, $\lambda(\mu),$ corresponding to the Jacobian matrix at $E_{1},$
@@
The statement of Proposition 3.10 says nothing about the criticality of the Hopf Bifurcation. Analytically, it is very difficult to compute the first Lyapunov coefficient of (3.1) at an equilibrium whose explicit expression we do not know. However, numerics suggest that the Hopf bifurcation is supercritical for $c=\sqrt{\frac{b}{1-{x^{*}}^{2}}}$ and subcritical for $c=-\sqrt{\frac{b}{1-{x^{*}}^{2}}}$.
4. Analysis of the singular case $c\to+\infty$ in Cases A and B
The aim of this section is the study of the dynamics of (3.1) when $|c|\to+\infty$. Many results concerning the asymptotic behaviour of the solution of the FN model as $|c|\to+\infty$ are available in the literature – see [21, 22, 23]. Our concern is to match these results with the local bifurcation study performed in Section 3. We are particularly interested in Cases A and B whose results have been described in Theorems 3.4 and 3.9. By changing the time scale of (3.1), taking into account $\tau=t/c$ and $\varepsilon=1/c^{2},$ we get an equivalent form for (3.1) as follows (written with respect to the slow time $\tau$):
$\left\{\begin{array}[]{l}\varepsilon\dot{x}=\displaystyle\left[y-\left(\frac{x^{3}}{3}-x\right)\right]=:f(x,y,\varepsilon)\\ \\ \dot{y}=\displaystyle-(x+by-a)=:g(x,y,\varepsilon)\end{array}\right$. (4.1)
[…]
A fast-slow system consists of differential equations where the rates of change of some variables are much greater than those of others. This leads to a two time-scale system. The general approach to this type of system starts by grouping the variables in two disjoint sets: fast variable $x$ and slow variable $y$.
By decomposing the dynamics into slow and fast phases, each can be studied independently, allowing a clearer understanding of its underlying behaviour. The parameter $\varepsilon$ introduces this separation in system (4.1).
@@ 4.1. The critical manifold and bifurcations
$\mathcal{C}_{0}=\{(x,y)\in\mathbb{R}^{2}:f(x,y,0)=0\}=\left\{(x,y)\in\mathbb{R}^{2}:y=\frac{x^{3}}{3}-x\right\}.$ (4.2)
@@
    The subset $S\subset\mathcal{C}_{0}$ is called normally hyperbolic provided that for all $e=(x_{0},y_{0})\in{S}$, we get:
    $\frac{\partial f}{\partial x}\left(e,0\right)\neq 0$
@@
    A normally hyperbolic set $S$ is called attracting (resp. repelling) when, for every $e\in{S}$, all eigenvalues of $\frac{\partial f}{\partial x}\left(e,0\right)$ possess negative (resp. positive) real part, respectively.
@@
With respect to (4.1) with critical manifold $\mathcal{C}_{0}$ (see (4.2)), the point $e\in\mathcal{C}_{0}$ is called a fold point if:
$\frac{\partial f}{\partial x}\left(e,0\right)=0\,,\quad\frac{\partial^{2}f}{\partial x^{2}}\left(e,0\right)\neq 0\quad\text{ and }\quad\frac{\partial f}{\partial y}\left(e,0\right)\neq 0.$
If $f\left(e,0\right)\neq 0$, the fold point is termed regular.
Coming back to system (4.1), we get $f\left(x,y,0\right)=x-\frac{x^{3}}{3}+y$ and, for $e=(x_{0},y_{0})\in\mathcal{C}_{0}$, one has:
$\frac{\partial f}{\partial x}\left(e,0\right)=1-x_{0}^{2}\neq 0\qquad\text{ if }\,\,x_{0}\neq\pm 1.$
Therefore, if $b=3/2$, the equilibria $E_{2}=\left(-1,2/3\right)$ and $E_{3}=(1,-2/3)$ are fold points (according to the Definition 4.2) since:
@@
and $S_{0}=\mathcal{C}_{0}\setminus\left\{E_{2},E_{3}\right\}$ is normally hyperbolic according to Definition 4.1. We may split $S_{0}$ into three disjoint open subsets:
@@
where $\mathcal{C}_{0\text{L}}$ and $\mathcal{C}_{0\text{R}}$ are attracting and $\mathcal{C}_{0\text{M}}$ is repelling. As a result, trajectories of (4.1) in the fast flow move towards either to $\mathcal{C}_{0\text{L}}$ or $\mathcal{C}_{0\text{R}}$ and backwards to $\mathcal{C}_{0\text{M}}$ – see Figure 5.
@@
Yet in the singular case ($\varepsilon=0$), depending on the values of $a,b\in\mathbb{R}$ in system (4.1), one, two, or three equilibria might arise as a consequence of the cubic curve’s intersection, $f(x,y,0)=0,$ and the line defined by xa+by=0x-a+by=0 ($\Leftrightarrow\dot{y}=0).$ If the system possesses three equilibria, only one of the following cases arise (cf. [10]):
@@
each of the three equilibria lies in $\mathcal{C}_{0\text{M}}$ or else
@@
each of the equilibria belongs to a different region $\mathcal{C}_{0\text{L}}$, $\mathcal{C}_{0\text{M}}$, and $\mathcal{C}_{0\text{R}}$.
When $\varepsilon=0,$ the independence of the time scales simplifies the system’s analysis. Fenichel’s Theorem [6] says that, near the normally hyperbolic part of $\mathcal{C}_{0}$, for $0<\varepsilon\ll 1,$ the perturbed system exhibits dynamics “analogous” to the singular system (in the sense of differentiable conjugacy), with a deviation of the order $\mathcal{O}\left(\varepsilon\right).$ Decreasing $\varepsilon$ makes the system trajectories increasingly similar to those of the singular system ($\varepsilon=0$). According to Fenichel’s Theorem, it is ensured that when $0<\varepsilon\ll 1,$ there is a set $S_{\varepsilon}$, $\mathcal{O}\left(\varepsilon\right)$–close to $S_{0}$ (in the $\mathbb{C}^{1}$–topology), which is locally hyperbolic and attracting (if $S_{0}$ is normally hyperbolic and attracting).
@@
The set $S_{\varepsilon}$ is known as the slow manifold for (4.1).
When the equilibrium of the system is a fold point, a subcritical (with respect to $b$) Hopf bifurcation arises, leading to the emergence of a periodic solution. In this section, the canard phenomenon described in [12, 13] links the local periodic solution generated by the Hopf bifurcation to the global limit cycle, like we proceed to explain (the canard point is correlated with the passage of a reduced problem equilibrium over the fold point under parameter variation – see Definition 4.6).
@@
A trajectory segment of a fast-slow system (4.1) is a canard if it stays within a $\mathcal{O}\left(\varepsilon\right)$–distance to a repelling branch of a slow manifold for a time that is $\mathcal{O}\left(1\right)$ on the slow time scale $\tau=t\varepsilon$.
Canards are related to equilibria of (4.1) that arise at fold points of the slow manifold. The next definitions deal with the general equation parametrised by $\lambda\in\mathbb{R}$:
@@
We now define a singular fold point, a different concept from that of a fold point; it is necessary to coincide with an equilibrium point of (4.1):
@@
With respect to (4.4), let $e=\left(x,y\right)$ denote a fold point. We refer to $e$ as a singular fold provided:
$f\left(e,\lambda,0\right)=0,$ $\frac{\partial f}{\partial x}\left(e,\lambda,0\right)=0,$ (4.5)
$\frac{\partial^{2}f}{\partial x^{2}}\left(e,\lambda,0\right)\neq 0,$ $\frac{\partial f}{\partial y}\left(e,\lambda,0\right)\neq 0\quad\text{and}\quad g\left(e,\lambda,0\right)=0.$
@@
A singular fold point $e$ (also known as a canard point) is called regular if:
$\frac{\partial g}{\partial x}\left(e,\lambda,0\right)\neq 0\quad\text{ and }\quad\frac{\partial g}{\partial\lambda}\left(e,\lambda,0\right)\neq 0.$ (4.6)
The dynamics near a regular singular fold point is explored in Theorem 4.7. The occurrence of canards is linked to a Hopf bifurcation, known in these systems as a singular Hopf bifurcation.
@@
Consider a fast-slow system of the form (4.4) where $\left(x,y\right)=\left(0,0\right)$ is canard point for $\lambda=0$, and
@@
Suppose that, when $\varepsilon=0$, a slow trajectory connects attracting and repelling regions of the critical manifold $\mathcal{C}_{0}$. Therefore, there is $\varepsilon_{0}>0$ and $\lambda_{0}>0$ so that for $0<\varepsilon<\varepsilon_{0}$ and $|\lambda|<\lambda_{0}$, the system possesses an equilibrium point $p\in\mathbb{R}^{2}$ around $(0,0)$ where $p$ approaches $(0,0)$ as $\left(\lambda,\varepsilon\right)$ tends to $(0,0)$. Moreover, there exists a smooth map $\lambda_{c}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that assigns every $\varepsilon\in\left(0,\varepsilon_{0}\right]$ to a real value $\lambda$, leading to a family of canards, defined asymptotically as
$\lambda_{c}\left(\sqrt{\varepsilon}\right)=-\varepsilon\left(A+B\right)+\mathcal{O}\left(\varepsilon^{3/2}\right).$
and In addition, there is a continuous map $\lambda_{H}:\left[0,\varepsilon_{0}\right]\to\mathbb{R}$ that associates each $\varepsilon\in\left[0,\varepsilon_{0}\right]$ to a value $\lambda$, associated to a Hopf bifurcations defined asymptotically as
$\lambda_{H}\left(\sqrt{\varepsilon}\right)=-\varepsilon B\varepsilon+\mathcal{O}\left(\varepsilon^{3/2}\right),$